(i) Primes 7 and 11 give 8 pairs, P = 8/36 = 2/9. (ii) P(different) = 1 – P(same) = 1 – 13/36 = 23/36. (iii) Favourable outcomes {HHT, HTH}, P = 2/8 = 1/4. (iv) Last digit 2 or 4, P = ...
Home/ganita manjari class 9 exercise 7.4 solutions
Discussion Forum Latest Questions
Virat
Asked: In: Class 9 Maths
(i) The sample space is S = {(Samosa, Chai), (Samosa, Lassi), (Pakora, Chai), (Pakora, Lassi), (Bhaji, Chai), (Bhaji, Lassi)}, giving 6 outcomes. (ii) The event selecting Samosa is E = {(Samosa, Chai), (Samosa, Lassi)}.
ganita manjari class 9 exercise 7.2 solutionsganita manjari class 9 exercise 7.3 solutionsganita manjari class 9 exercise 7.4 solutionsncert class 9 ganita manjari chapter 7 question answersample space of all possible snack and drink combinationsselecting samosa as a snackthere are 3 popular snacks availablevillage fair
Virat
Asked: In: Class 9 Maths
(i) S = {1H, 1T, 2H, 2T, 3H, 3T, 4H, 4T, 5H, 5T, 6H, 6T} with 12 outcomes. (ii) S = {-4, -3, -2, -1, 0, 1, 2, 3, 4}. (iii) By color, S = {Green, Red}; by distinct balls, S = ...
box containing 5 green and 7 red ballschapter 9 the mathematics of maybe: introduction to probability question answerganita manjari class 9 exercise 7.3 solutionsganita manjari class 9 exercise 7.4 solutionsncert solutions – the mathematics of maybe: introduction to probabilityrolling a die and tossing a coin together